please show work!!!!

Part 2, Instructions: Read the information below and use this to answer question 9-11 on page 8 Section 9. OVERTIME Pursuant to the provisions of the Fair Labor Standards Act, non-exempt employees shall be paid overtime at their appropriate rate of not less than one and one-half(1 %) times their rate of pay for each hour worked in a work period in excess of the maximum hours allowable, which are applicable to the type of employment in which the employee is engaged 9.1 92 Pursuant to the specific provisions of 29 U.S.C., Section 207 (k) any non-exempt employee engaged in law enforcement activities shall be paid overtime at the special hourly base rate for all hours worked in excess of eighty- five and a half (85.5) hours in a fourteen (14) day work period. (a) any non-exempt employce engaged in fire protection activities shall be paid (b) overtime at the special hourly base rate for all hours worked in excess of two hundred and twelve (212) hours in a twenty-eight (28) day work period. any non-exempt employee not covered by Section 9.2 (a) and (b), who receives state supplemental pay and/or city millage, shall be paid overtime worked in excess of forty (40) hours in a work period. (c) 9.3 Pursuant to the specific provisions of 29 US.С., Section 207 (a), any non-exempt employee shall be paid overtime at the regular hourly base rate for all hours worked itn excess of forty (40) hours in a seven (7) consecutive day work period. Annual leave, sick leave, civil leave, military leave, maternity leave, as well as all other forms of leave, and official and unofficial city holidays will not be counted as hours worked toward the computation of overtime at the completion of the maximum allowable hours in the respective work period. 9.4 If it becomes necessary for a non-exempt employee to work on an official city holiday as designated by the City Council, the appointing authority should adjust the employee's work schedule to allow another day off during that work period as a substitution for the holiday. If such a substitution is not possible, then for working on the holiday the employee shall be paid double time. Under no circumstances shall an employee receive pay less than an amount equal to straight time for all hours worked plus holiday pay as established in Rule X, section 1.2. However, on New Year's Day, Mardi Gras, Independence Day, Labor Day, Thanksgiving Day, Christmas Day or on their days of observance as officially designated by the City Council, employees who are required to work shall be paid double time and one-half Under no circumstances shall an employee receive less than an amount equal to time and one-half for all hours worked, plus holiday pay as established in Rule X, Section 1.2 9.5
Questions for Section 3, Part 2 begin here 9. John makes $10.00 an hour normally. He has to work overtime hours on Thanksg How much will he make per hour for working that day? A. S10.00 B. $15.00 C. $20.00 D. $25.00 Mary is not in law enforcement or fire protection and usualy works 35 hours a week (5- 7-hour days). She had to stay late 2 's hours on Monday, called in sick for 7 hours on Tuesday, and worked on a special project for 3 hours on Saturday. How many hours of overtime has Mary earned? Α.0 B. 1 C. 2 D. 3 E. 6 Assume that Mary (from question 10) did not call in sick on Tuesday. Now, how many hours of overtime (paid at time and a half) has Mary earned? 11. A. 0 B. C. 2% D. 3 E. 6
Answer #1

Answer 9

Normal hourly rate is $10,

As per clause in 9.1 minimum overtime rate shall be 1.5 times, hence his overtime wage will be 10*1.5 = $15

Hence option B

Answer 10

If mary is not in law enforcement and fire protection clause 9.2 (c) will be applicable. Hence her weekly working hour is 40 hours after which overtime will be applicable.

In given week she usually work for 35 hours.

Extra Work was on monday for 2.5 hours and saturday for 3.5 hours on saturday. Total extra work = 2.5+3.5 = 6 hours.

However she was sick and took off on tuesday for 7 hours.

Hence net extra work = 6-7 = -1

Hence she is not entitled for overtime.

Hence option A is correct

Answer 11

If Mary didnot take 7 hour then she worked for total of 35 (normal) +6 (extra) = 41 hours.

Minimum requirement is 40 hours.

Hence she is entitled for 41-40= 1 hour overtime

Hence option B is correct